The King Food Corp. currently has no debt in its capital structure. The beta of its capital is 1.5. The King Food Corp’s free cash flow is expected to equal £30 million next year. This cash flow is expected to grow at 2% per year for the foreseeable future. King Food Corp. is considering changing its capital structure by issuing debt and using the proceeds to buy back stock. It will do so in such a way that it will have a 50% debt-to-equity ratio (D/E=50%) after the change, and it will maintain this debt-equity ratio forever. Assuming King Food Corp’s pretax cost of debt will be 5%. King Food Corp. faces a corporate tax rate 40%. Assuming that the CAPM holds, the risk-free rate is 3%, and the expected market index risk premium is 8%.

                

(b) Using the information provided and your calculations in part (a) and (b), determine the value of tax shield acquired by King Food Corp. if it changes its capital structure with a 50% debt-to-equity ratio (D/E=50%).                                   


1
Expert's answer
2022-03-03T14:58:59-0500

a)the cost of equity before the change in capital structure:

"CARM=3+1.5(8-3)=10.5"


after change in the capital structure:

"WACC=0.5\times5(1-0.4)+0.5\times10.5=1.5+5.25=6.75"